The average peak current in a lightning strike is 15 kA. What is the magnitude of the magnetic flux density at a distance of 10 meters from the flash channel? What about 100 meters from the flash channel?



1
Expert's answer
2021-08-29T16:56:49-0400

"B_i=\frac{\mu_0I}{2\pi r_i},"

"B_1=3\cdot 10^{-4}~T,"

"B_1=3\cdot 10^{-5}~T."